# Detection of a Type-C QPO during the soft-to-hard transition in Swift J1727.8 –1613

### v1

@@ Abstract
Context. Timing analysis of accreting systems is key to probe the structure and dynamics around compact objects. In Black-Hole Low-Mass X-ray Binaries (BH LMXBs), the compact object accretes matter from a low-mass companion star via Roche Lobe overflow, forming an accretion disk, and occasionally exhibiting bright eruptions. The BH LMXB Swift J1727.8–1613 (hereafter J1727), recently underwent one of the brightest outbursts ever recorded in X-rays, in August 2023.
Aims. This analysis aims to study the timing properties of J1727, in the decaying phase of its outburst, using high-time resolution XMM–Newton data.
Methods. We analyzed J1727’s power spectrum (PS) and cross spectrum (CS), which we modeled with Lorentzians. The PS reveals how the source’s power is distributed across frequencies, and the Real and Imaginary parts of the CS compare the displacement of the light curves in different energy bands across the observations. Finally, we simultaneously derived the phase lags and the coherence, using a constant phase lag model.
Results. While the first (soft-state) observation does not show any strong variability, the two harder observations exhibit quasi-periodic oscillations (QPOs). Because the QPO is more significantly detected in the Imaginary part of the CS than in the PS, we refer to it as the ”Imaginary QPO”. The QPO is more prominent in the soft 0.3–2 keV band than in the hard 2–12 keV band. As the source evolves towards the hard state, the Imaginary QPO shifts to lower frequencies, the broadband fractional rms amplitude in the 0.3–2 keV energy band increases, while the rms covariance of the Imaginary QPO decreases. Simultaneously, the phase lags increase and the coherence function drops at the Imaginary QPO frequency.
Conclusions. In the elusive soft-to-hard transition of J1727, the first XMM-Newton observations of the source reveal an Imaginary QPO also detected in the PS, exhibiting the properties of a type-C QPO.
@@ 1 Introduction
Black hole Low-Mass X-ray Binaries (BH LMXBs) consist of a black hole accreting matter from a low-mass companion star through Roche Lobe overflow Frank et al. (2002). Several of these sources display outbursts, during which their flux increases by a few orders of magnitude over a few weeks to months (Zdziarski et al., 2004; Done et al., 2007; Hirsch et al., 2020). The spectral and timing properties of these sources exhibit dramatic changes throughout the outburst evolution, transitioning from the Hard state (HS) to the Soft state (SS; Remillard & McClintock, 2006). In the HS, the emission from the hot Comptonized medium (often referred to as the ”corona”) dominates the X-ray spectrum. In contrast, during the SS, a multi-blackbody emission from the accretion disk dominates the high-energy spectrum (see Shakura & Sunyaev, 1973). The transition between these two spectral states is often referred to as the hard-intermediate state (HIMS) and the soft-intermediate state (SIMS; Belloni et al., 2005), which are less commonly observed than the HS and the SS, because of their short lifespan (Dunn et al., 2010). Swift J1727.8–1613 (hereafter J1727), is one of the few sources observed in the decaying phase of the outburst. First detected on August 24, 2023 (Dichiara et al., 2023; Negoro et al., 2023), the source’s emission peaked at around 7 Crabs in the 2–20 keV energy band, exhibiting an exceptionally bright outburst. J1727 is the first BH LMXB observed with the Imaging X-ray Polarimetry Explorer (IXPE; Weisskopf et al., 2022) across the entire outburst – from the brightening in the hard state (Veledina et al., 2023), through the hard–to–soft transition (Ingram et al., 2024), to the soft state (Svoboda et al., 2024), and finally in the dim hard state (Podgorný et al., 2024). Optical spectroscopy during quiescence suggests the presence of a black hole primary (Mata Sánchez et al., 2024), located at $5.5^{1.4}_{1.1}$ kpc (Burridge et al., 2025), defining the system as a Low-Mass X-ray Binary (LMXB; Castro-Tirado et al., 2023).
@@
Throughout the outburst – traced by a ”q” shape in the hardness intensity diagram (HID; e.g. Homan et al., 2001; Fender et al., 2004) – transient BHXBs exhibit fast X-ray variability. Quasi-periodic oscillations (QPOs; van der Klis, 1989; Psaltis et al., 1999; Belloni et al., 2002; Giannios & Spruit, 2004) are among the components that can be tracked during an outburst, providing insight into the inner parts of the system. Low-Frequency QPOs (LFQPOs) are the most commonly observed, with frequencies ranging from $\sim$0.01 Hz to $\sim$30 Hz (Norris & Wood, 1987; Wijnands & van der Klis, 1998; Casella et al., 2004; Belloni et al., 2005). Based on the strength of the broadband variability and QPO characteristics, LFQPOs can be classified into three main types: Type-A, -B and -C (Casella et al., 2005).
@@
Type-C QPOs are characterized by a strong and narrow peak in the power spectra (PS). The observed fractional root mean square (rms) amplitude ranges from 15% to 30%, and the QPO appears simultaneously with strong broadband variability (Casella et al., 2004; Belloni et al., 2005). The origin of these QPOs remains an open question. Nevertheless, the leading suspects proposed in different models are instabilities in the accretion flow (Tagger & Pellat, 1999; Titarchuk & Fiorito, 2004; Mondal et al., 2009), geometrical effects due to the Lense–Thirring precession (Stella & Vietri, 1998; Stella et al., 1999; Ingram et al., 2009; Kubota et al., 2024), time-dependent Comptonization (García et al., 2021; Mastichiadis et al., 2022; Bellavita et al., 2022), and a precessing jet (Ma et al., 2021). Type-C QPOs were detected in the HIMS of J1727 with different X-ray telescopes such as the Neil Gehrels Swift Observatory (Palmer & Parsotan, 2023), the Neutron Star Interior Composition Explorer (NICER; Draghis et al., 2023; Bollemeijer et al., 2023; Liao et al., 2024; Rawat et al., 2025), Insight-Hard X-ray Modulation Telescope (Insight-HXMT; Yu et al., 2024; Chatterjee et al., 2024) and the Imaging X-ray Polarimetry Explorer (IXPE; Zhao et al., 2024; Ingram et al., 2024).
@@
Probing the QPO’s characteristics during a state transition is crucial to understand their origin and formation. Méndez et al. (2024) simultaneously fitted the PS and the Real and Imaginary parts of the cross spectra (CS) using a multi-Lorentzian model in different LMXBs: GX 339–4, GRS 1915+105, and during the soft–to–hard transition of MAXI J1820+070. Their analyses show the necessity to fit the Real and Imaginary parts of the CS, as some QPOs can be significant in the CS but not in the PS. In the soft energy band (below 1 keV), König et al. (2024) and Fogantini et al. (2025) performed a similar analysis on the High-mass X-ray binary (HMXB) Cyg X-1 using NICER data. Despite the differences in the companion’s mass, several of these sources (including J1727), exhibit a sudden jump in the phase lags and a coherence drop at a specific frequency, which can be attributed to an additional narrow Lorentzian visible only in the Imaginary part of the CS. This narrow Lorentzian is likely induced by the variability from the corona. Studies in the rising and decaying phases of the outburst reveal these features only arising during the decaying phase of the soft–to–hard transition (Alabarta et al., 2022; Ma et al., 2023; Bellavita et al., 2025), suggesting the disk inner radius expansion, or feedback from the corona as their origin.
This work presents the first timing analysis of J1727 in the soft–to–hard transition with XMM–Newton data, using the multi-Lorentzian decomposition of the PS and CS framework presented in Méndez et al. (2024). From the CS of two different energy bands, we derived the phase lags, which trace the phase angle in the complex Fourier plane of the cross vector for correlated signals (van der Klis et al., 1987; van der Klis, 1989; Vaughan et al., 1997; Nowak et al., 1999a). The coherence function – quantifying the linear correlation between the two signals depending on the Fourier frequency (Vaughan & Nowak, 1997) – is then calculated.
@@
This paper is structured as follows: we present the data set in Section 2, the multi-Lorentzian model fitting of the PS and CS is presented in Section 3, and discussed in Section 4, alongside the phase lags and coherence fitting.
@@ 2 Data Sets
We observed J1727 with the XMM–Newton’s European Photon Imaging Camera (EPIC)-PN spectrograph (covering the 0.15 – 12 keV energy band; Jansen et al., 2001; Strüder et al., 2001) during the soft–to–hard transition (see Table 2, PID 88500; Castro Segura, 2020). The data were processed using the Science Analysis Software (SAS; Gabriel et al., 2004) pipeline from ESA. The calibrated photon event files for the PN camera were produced with epproc. We extracted, filtered for pile-up, and rebinned (0.1 s) the light curves using evselect, for all three observations. We defined Good Time Intervals (GTIs) to remove the telemetry gaps (see Fig. 4), and computed the power density and cross spectra presented in section 3.
@@ 3 Analysis and results
Table 1: Best fitting parameters of the PS and CS for both observations in the soft–to–hard transition. “Comp” refers to the model components and L1-3 refer to the three fitted Lorentzians (defined in equations 1 for the PS and in equation 3 for the CS).
[…]
@@
[…]
@@
Figure 1: Left panel: Power spectra of the second (top) and third (bottom) EPIC-PN observations of Swift J1727.8–1613 in the 0.3–2 keV (black) and 2–12 keV (orange) energy bands. Right panel: Real (green) and Imaginary (pink) parts of the cross spectra of the same two observations using the 0.3–2 keV and 2–12 keV energy bands, rotated by 45°(see Section 3.2 for more details). The vertical red line shows the position of the central frequency $\nu_{0}$ of the Imaginary QPO. The total model is shown with the continuous line and the individual Lorentzians are in dotted lines. Below each fit are the residuals defined as $\mathrm{\Delta\chi=}$ (data–model)/error.
Using the General High-Energy Aperiodic Timing Software (GHATS [^1] http://astrosat-ssc.iucaa.in/uploads/ghats_home.html), we calculated the PS in the 0.3–2 keV and 2–12 keV energy bands, and the cross spectra between those same two bands. We fixed the sampling time to 0.006 s (200×200$\times$200 × the native instrumental resolution), which defines the Nyquist frequency at $\mathrm{0.5/0.006=83.3}$ Hz. Based on the typical length of the GTI segments allowed by the telemetry gaps, we set the interval length for computing the PS to 2048× $\mathrm{2048\ \times\ 0.006s\approx 12.3\ s}$, corresponding to a minimum frequency of $\sim$0.08 Hz. Consequently, the second observation (ObsID 0885000701) ended up with 641 segments against 1605 segments for the third observation (ObsID 0885000801). To increase the Signal to Noise Ratio at high frequencies, we logarithmically rebinned the PS and CS, by a factor of $\mathrm{10^{0.01}\approx 1.023}$. The PS and CS were normalized to fractional rms units (van der Klis, 1989). Poisson noise and cross-talk levels were estimated empirically from the 30–80 Hz frequency range, where no signal is detected, and subtracted from the PS and the Real part of the CS. Therefore, the PS and CS are fitted only up to 30 Hz. Using the PS and CS, we computed the phase lags and the coherence for all observations, then fitted with the X-Ray Spectral Fitting Package (Xspec; Arnaud, 1996).
@@ 3.1 The power spectra
Unsurprisingly, no significant variability was observed in the PS of the first (soft state) observation. Therefore, this analysis focuses on the last two observations during the soft–to–hard transition. The PS were computed using GHATS and fitted in Xspec with an additive multi-Lorentzian model for each observation (see Fig. 1). Following Méndez et al. (2024), the Lorentzians are assumed to be coherent between the two energy bands across the Fourier frequencies, but incoherent between each other. Thus, the PS in two energy bands, x and y, can be defined as a sum of Lorentzians:
@@
[…]
@@
Table 1 gives the best-fitting parameters from the Xspec fitting with three Lorentzians for each observation. The central frequency $\mathrm{\nu_{0}}$ and the FWHM of the Lorentzians in the 2–12 keV energy band are linked to the values from the low energy band (0.3–2 keV). Only the fractional rms amplitude remains a free parameter in both bands. The quality factor $Q$ is defined as the ratio between the central frequency of the QPO and the FWHM. The best fits using the multi-Lorentzian model give $\mathrm{\chi^{2}\ /\ dof\sim 1.3}$ and $\mathrm{\chi^{2}\ /\ dof\sim 1.9}$, for the second and third observation, respectively. The model describes the data correctly in both cases, without prominent features in the residuals. Overall, the PS has a greater rms amplitude at higher energy, with a total rms of $\sim$17in 0.3–2 keV versus $\sim$30in 2–12 keV.
Additionally, the rms of the first and last Lorentzians (referred to as L1 and L3) increase in the 0.3–2 keV energy band as the outburst evolves from the second to the third observation: the rms goes from $\sim$16.4to $\sim$17.1for L1 and from $\sim$4.5to $\sim$6.4for L3. In the 2–12 keV energy band, the rms similarly increases for L3 but decreases for L1, from $\sim$6.3to $\sim$3.7%. On the other hand, the rms amplitude of L2 decreases in both energy bands: from $\sim$3.5to $\sim$2.9in the 0.3–2 keV energy band and from $\sim$6.3to $\sim$3.7in the 2–12 keV energy band.
Furthermore, the central frequencies of the three Lorentzians shift toward lower frequencies as the source hardens, decreasing by $\sim$16%, $\sim$23%, and $\sim$28for L1, L2, and L3, respectively, between the second and third observation. The Lorentzians also become narrower, with a decrease in the FWHMs of $\sim$18%, $\sim$62and $\sim$20between the second and third observation (see Table 1). As a result, as the source hardens, the quality factor of the second Lorentzian increases from 2.07 to 4.12, but decreases for L3 from 0.60 to 0.54. Parallelly, the $Q$ factor of L1 remains constant around 0.23 within the errors.

### v2

# Detection of a type-C quasi-periodic oscillation during the soft-to-hard transition in Swift J1727.8 –1613

@@ Abstract
Context. Timing analysis of accreting systems is key to probing the structure and dynamics around compact objects. In black-hole low-mass X-ray binaries (BH LMXBs), the compact object accretes matter from a low-mass companion star via Roche-Lobe overflow and forms an accretion disk that occasionally exhibits bright eruptions. The BH LMXB Swift J1727.8–1613 (hereafter J1727) underwent one of the brightest outbursts ever recorded in X-rays in August 2023.
Aims. We study the timing properties of J1727 in the decaying phase of its outburst based on XMM–Newton data with a high-time resolution.
Methods. We analyzed the power spectrum (PS) and cross spectrum (CS) of J1727, which we modeled with Lorentzians. The PS reveals the power distribution of the source across frequencies, and the real and imaginary parts of the CS compare the displacement of the light curves in different energy bands for the different observations. Finally, we simultaneously derived the phase lags and the coherence using a constant phase-lag model.
Results. While the first (soft-state) observation shows no strong variability, the two harder observations exhibit quasi-periodic oscillations (QPOs). Because the QPO is more significantly detected in the imaginary part of the CS than in the PS, we refer to it as the ”imaginary QPO”. The QPO is more prominent in the soft 0.3–2 keV band than in the hard 2–12 keV band. As the source evolves toward the hard state, the imaginary QPO shifts to lower frequencies, the broadband fractional rms amplitude in the 0.3–2 keV energy band increases, and the rms covariance of the imaginary QPO decreases. Simultaneously, the phase lags increase, and the coherence function drops at the imaginary QPO frequency.
Conclusions. This analysis provides the first type-C QPO detection in a BH XB during the soft–to–hard transition using XMM–Newton data. The QPO is detected at particularly low energy (0.3–2 keV). Notably, the QPO is significantly detected in the imaginary part of the CS and the PS. Thus, we confirm the physical origin of the coherence drop and the phase-lag excess, which were only observed with NICER before.
@@ 1 Introduction
Black hole low-mass X-ray binaries (BH LMXBs) consist of a black hole accreting matter from a low-mass companion star through Roche-Lobe overflow Frank et al. (2002). Several of these sources display outbursts, during which their flux increases by a few orders of magnitude over a few weeks to months (Zdziarski et al., 2004; Done et al., 2007; Hirsch et al., 2020). The spectral and timing properties of these sources change dramatically throughout the outburst evolution and transition from the Hard state (HS) to the Soft state (SS; Remillard & McClintock, 2006). In the HS, the emission from the hot Comptonized medium (often referred to as the ”corona”) dominates the X-ray spectrum. In contrast, during the SS, a multiblackbody emission from the accretion disk dominates the high-energy spectrum (see Shakura & Sunyaev, 1973). The transition between these two spectral states is often referred to as the hard-intermediate state (HIMS) and the soft-intermediate state (SIMS; Belloni et al., 2005). These states are far more rarely observed than the HS and the SS because their lifespan is short (Dunn et al., 2010). Swift J1727.8–1613 (hereafter J1727) is one of the few sources that were observed in the decaying phase of the outburst. First detected on August 24, 2023 (Dichiara et al., 2023; Negoro et al., 2023), the source emission peaked at around 7 Crab in the 2–20 keV energy band; this event was an exceptionally bright outburst. J1727 is the first BH LMXB that was observed with the Imaging X-ray Polarimetry Explorer (IXPE; Weisskopf et al., 2022) throughout the entire outburst, from the brightening in the hard state (Veledina et al., 2023) through the hard–to–soft transition (Ingram et al., 2024) to the soft state (Svoboda et al., 2024), and finally, in the dim hard state (Podgorný et al., 2024). Optical spectroscopy during quiescence suggests the presence of a black hole primary (Mata Sánchez et al., 2024), located at $5.5^{1.4}_{1.1}$ kpc (Burridge et al., 2025), which defines the system as a low-mass X-ray binary (LMXB; Castro-Tirado et al., 2023).
@@
Throughout the outburst, which is visible as a ”q” shape in the hardness-intensity diagram (HID; e.g. Homan et al., 2001; Fender et al., 2004), transient BH XBs show fast X-ray variability. Quasi-periodic oscillations (QPOs; van der Klis, 1989; Psaltis et al., 1999; Belloni et al., 2002; Giannios & Spruit, 2004) are among the components that can be tracked during an outburst. They provide insight into the inner parts of the system. Low-frequency QPOs (LFQPOs) are the most commonly observed events. Their frequencies range from $\approx$0.01 Hz to $\approx$30 Hz (Norris & Wood, 1987; Wijnands & van der Klis, 1998; Casella et al., 2004; Belloni et al., 2005). Based on the strength of the broadband variability and QPO characteristics, LFQPOs can be classified into three main types: type A, B and C (Casella et al., 2005).
@@
Type-C QPOs are characterized by a strong and narrow peak in the power spectra (PS). The observed fractional root mean square (rms) amplitude ranges from 15% to 30%, and the QPO is simultaneously observed with strong broadband variability (Casella et al., 2004; Belloni et al., 2005). The origin of these QPOs remains an open question. Nevertheless, the leading suspects proposed in different models are instabilities in the accretion flow (Tagger & Pellat, 1999; Titarchuk & Fiorito, 2004; Mondal et al., 2009), geometrical effects caused by the Lense–Thirring precession (Stella & Vietri, 1998; Stella et al., 1999; Ingram et al., 2009; Kubota et al., 2024), time-dependent Comptonization (García et al., 2021; Mastichiadis et al., 2022; Bellavita et al., 2022), and a precessing jet (Ma et al., 2021). Type-C QPOs were detected in the HIMS of J1727 with different X-ray telescopes, such as the Neil Gehrels Swift Observatory (Palmer & Parsotan, 2023), the Neutron Star Interior Composition Explorer (NICER; Draghis et al., 2023; Bollemeijer et al., 2023; Rawat et al., 2025; Liao et al., 2025), Insight-Hard X-ray Modulation Telescope (Insight-HXMT; Yu et al., 2024; Chatterjee et al., 2024), and the Imaging X-ray Polarimetry Explorer (IXPE; Zhao et al., 2024; Ingram et al., 2024).
@@
It is crucial to probe the QPO’s characteristics during a state transition to understand their origin and formation. Méndez et al. (2024) simultaneously fit the PS and the real and imaginary parts of the cross spectra (CS) using a multi-Lorentzian model in different LMXBs: GX 339–4, GRS 1915+105, and during the soft–to–hard transition of MAXI J1820+070. Their analyses showed that the real and imaginary parts of the CS must be fit because some QPOs can be significant in the CS, but not in the PS. In the soft-energy band (below 1 keV), König et al. (2024) and Fogantini et al. (2025) performed a similar analysis of the high-mass X-ray binary (HMXB) Cyg X-1 using NICER data. Despite the differences in the companion mass, several of these sources (including J1727) showed a sudden jump in the phase lags and a coherence drop at a specific frequency. This can be attributed to an additional narrow Lorentzian that is only visible in the imaginary part of the CS. This narrow Lorentzian is likely caused by the variability of the corona. Studies in the rising and decaying phases of the outburst revealed that these features only arise during the decaying phase of the soft–to–hard transition (Alabarta et al., 2022; Ma et al., 2023; Bellavita et al., 2025). This suggests that the inner radius of the disk expands, or that it might come from feedback processes from the corona. We present the first timing analysis of J1727 in the soft–to–hard transition with XMM–Newton data. We used the multi-Lorentzian decomposition of the PS and CS framework presented by Méndez et al. (2024). From the CS of two different energy bands, we derived the phase lags, which trace the phase angle in the complex Fourier plane of the cross vector for correlated signals (van der Klis et al., 1987; van der Klis, 1989; Vaughan et al., 1997; Nowak et al., 1999a). We then calculated the coherence function, which quantifies the linear correlation between the two signals depending on the Fourier frequency (Vaughan & Nowak, 1997).
@@
This paper is structured as follows: We present the data set in Section 2 and the multi-Lorentzian model fitting of the PS and CS in Section 3. We discuss these results in Section 4, together with the phase lags and coherence fitting.
@@ 2 Data Sets
We observed J1727 with the XMM–Newton European Photon Imaging Camera (EPIC)-PN spectrograph (covering the 0.15 – 12 keV energy band; Jansen et al., 2001; Strüder et al., 2001) during the soft–to–hard transition (see Table 2, PID 88500; Castro Segura, 2020). The data were processed using the Science Analysis Software (SAS; Gabriel et al., 2004) pipeline from ESA. The calibrated photon event files for the PN camera were produced with epproc. We extracted, filtered for pile-up, and rebinned (0.1 s) the light curves using evselect for all three observations. We defined good time intervals (GTIs) to remove the telemetry gaps (see Fig. 4), and we computed the power density and cross spectra, which we present in section 3.
@@ 3 Analysis and results
Table 1: Best-fitting parameters of the PS and CS for the observations in the soft–to–hard transition. 111“Comp” refers to the model components and L1-3 refer to the three fit Lorentzians (defined in equations 1 for the PS and in equation 3 for the CS).
Epoch Comp. $\mathrm{\nu_{0}}$ [Hz] FWHM [Hz] $Q$ $\mathrm{=\frac{\nu_{0}}{FWHM}}$ $\mathrm{rms_{PS,\ 0.3-2keV}}$ [%] $\mathrm{rms_{PS,\ 2-12keV}}$ [%] Phase lags [rad]
L1 0.26 $\mathrm{\pm}$ 0.01 1.13 $\mathrm{\pm}$ 0.02 0.23 $\mathrm{\pm}$ 0.01 16.38 $\mathrm{\pm}$ 0.08 25.5 $\mathrm{\pm}$ 0.2 0.010 $\mathrm{\pm}$ 0.005
XMM2 L2 2.59 $\mathrm{\pm}$ 0.04 1.25 $\mathrm{\pm}$ 0.14 2.07 $\mathrm{\pm}$ 0.23 3.5 $\mathrm{\pm}$ 0.2 6.3 $\mathrm{\pm}$ 0.7 0.70 $\mathrm{\pm}$ 0.05
L3 3.48 $\mathrm{\pm}$ 0.24 5.75 $\mathrm{\pm}$ 0.30 0.60 $\mathrm{\pm}$ 0.05 4.5 $\mathrm{\pm}$ 0.2 15.0 $\mathrm{\pm}$ 0.6 0.26 $\mathrm{\pm}$ 0.03
@@
L1 0.218 $\mathrm{\pm}$ 0.006 0.93 $\mathrm{\pm}$ 0.01 0.234 $\mathrm{\pm}$ 0.007 17.13 $\mathrm{\pm}$ 0.07 24.6 $\mathrm{\pm}$ 0.1 0.016 $\mathrm{\pm}$ 0.003
XMM3 L2 1.98 $\mathrm{\pm}$ 0.01 0.48 $\mathrm{\pm}$ 0.05 4.12 $\mathrm{\pm}$ 0.43 2.9 $\mathrm{\pm}$ 0.1 3.7 $\mathrm{\pm}$ 0.3 0.81 $\mathrm{\pm}$ 0.04
L3 2.48 $\mathrm{\pm}$ 0.08 4.6 $\mathrm{\pm}$ 0.1 0.54 $\mathrm{\pm}$ 0.02 6.4 $\mathrm{\pm}$ 1.4 17.3 $\mathrm{\pm}$ 0.2 0.32 $\mathrm{\pm}$ 0.01
@@
Figure 1: Left panel: Power spectra of the second (top) and third (bottom) EPIC-PN observations of Swift J1727.8–1613 in the 0.3–2 keV (black) and 2–12 keV (orange) energy bands. Right panel: real (green) and imaginary (pink) parts of the cross spectra of the same two observations using the 0.3–2 keV and 2–12 keV energy bands, rotated by 45°(see Sect. 3.2 for more details). The vertical red line shows the position of the central frequency $\nu_{0}$ of the imaginary QPO. The total model is shown with the continuous line, and the individual Lorentzians are shown as dotted lines. The residuals defined as $\mathrm{\Delta\chi=}$ (data–model)/error are plotted below each fit.
Using the General High-Energy Aperiodic Timing Software (GHATS [^2] http://astrosat-ssc.iucaa.in/uploads/ghats_home.html), we calculated the PS in the 0.3–2 keV and 2–12 keV energy bands and the cross spectra between those same two bands. We fixed the sampling time to 0.006 s ($200\times$ the native instrumental resolution), which defines the Nyquist frequency at $\mathrm{0.5/0.006=83.3}$ Hz. Based on the typical length of the GTI segments allowed by the telemetry gaps, we set the interval length for computing the PS to $\mathrm{2048\ \times\ 0.006s\approx 12.3\ s}$, which corresponds to a minimum frequency of $\approx$0.08 Hz. Consequently, the second observation (ObsID 0885000701) had 641 segments, and the third observation (ObsID 0885000801) had 1605 segments. To increase the signal-to-noise ratio at high frequencies, we logarithmically rebinned the PS and CS by a factor of $\mathrm{10^{0.01}\approx 1.023}$. The PS and CS were normalized to fractional rms units (van der Klis, 1989). The Poisson noise and cross-talk levels were estimated empirically from the 30–80 Hz frequency range, where no signal was detected, and they were subtracted from the PS and the real part of the CS. Therefore, the PS and CS were only fit up to 30 Hz. Using the PS and CS, we computed the phase lags and the coherence for all observations, and we then fit them with the X-Ray Spectral Fitting Package (Xspec; Arnaud, 1996).
@@ 3.1 The power spectra
No significant variability was observed in the PS of the first (soft state) observation, as expected. This analysis, therefore, focuses on the last two observations during the soft–to–hard transition. The PS were computed using GHATS and were fit in Xspec with an additive multi-Lorentzian model for each observation (see Fig. 1). Following Méndez et al. (2024), the Lorentzians were assumed to be coherent between the two energy bands across the Fourier frequencies, but incoherent between each other. Thus, the PS in two energy bands, x and y, can be defined as a sum of Lorentzians,
@@
with $\mathrm{\textit{i}=1,\ 2,\ 3}$ for the three Lorentzians, $A_{i}$, $B_{i}$ are the integrated power from zero to infinity of the Lorentzians in the band x and y, respectively, $\mathrm{\nu_{0;\textit{i}}}$ is the central frequency for each individual Lorentzian, and $\mathrm{\Delta_{\textit{i}}}$ is their full width at half maximum (FWHM). Since for each Lorentzian function we assumed the coherence function to be $\mathrm{\gamma^{2}_{\textit{xy;i}}(\nu)=1}$ and the modulus square of the CS to be $\mathrm{|G_{\textit{xy}}(\nu)|^{2}=}$ $A_{i}$ $B_{i}$ $\mathrm{L^{2}(\nu;\nu_{0;\textit{i}},\Delta_{\textit{i}})}$, the former is defined as
@@
Table 1 gives the best-fitting parameters from the Xspec fitting with three Lorentzians for each observation. The central frequency $\mathrm{\nu_{0}}$ and the FWHM of the Lorentzians in the 2–12 keV energy band are linked to the values from the low-energy band (0.3–2 keV). Only the fractional rms amplitude remains a free parameter in both bands. The quality factor $Q$ is defined as the ratio of the central frequency of the QPO and the FWHM. The best-fits using the multi-Lorentzian model give $\mathrm{\chi^{2}\ /\ dof\approx 1.3}$ and $\mathrm{\chi^{2}\ /\ dof\approx 1.9}$ for the second and third observation, respectively. The model describes the data correctly in both cases, without prominent features in the residuals. The PS has a greater rms amplitude at higher energy overall, with a total rms of $\approx$17% in 0.3–2 keV versus $\approx$30% in 2–12 keV.
Additionally, the rms of the first and last Lorentzians (referred to as L1 and L3) increase in the 0.3–2 keV energy band as the outburst evolved from the second to the third observation: the rms increases from $\approx$16.4% to $\approx$17.1% for L1 and from $\approx$4.5% to $\approx$6.4% for L3. In the 2–12 keV energy band, the rms similarly increases for L3, but decreases for L1 from $\approx$6.3% to $\approx$3.7%. On the other hand, the rms amplitude of L2 decreases in both energy bands from $\approx$3.5% to $\approx$2.9% in the 0.3–2 keV energy band and from $\approx$6.3% to $\approx$3.7% in the 2–12 keV energy band.
Furthermore, the central frequencies of the three Lorentzians shift toward lower frequencies as the source hardens. They decrease by $\approx$16%, $\approx$23%, and $\approx$28% for L1, L2, and L3, respectively, between the second and third observation. The Lorentzians also become narrower, with a decrease in the FWHMs of $\approx$18%, $\approx$62% and $\approx$20% between the second and third observation (see Table 1). As a result, when the source hardens, the quality factor of the second Lorentzian increases from 2.07 to 4.12, but decreases for L3 from 0.60 to 0.54. The $Q$ factor of L1 simultaneously remains constant around 0.23 within the errors.